OpenStudy (anonymous):

A polyhedron has 10 faces and 24 edges. What is the number of vertices of the polyhedron? 12 16 10 14

OpenStudy (anonymous):

do u know Euler's formula?

OpenStudy (anonymous):

F+V-E=2

OpenStudy (anonymous):

@julie001 F+V=E+2?

OpenStudy (anonymous):

whcih one is it?

OpenStudy (anonymous):

mine

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

can u do it from there??

OpenStudy (anonymous):

10+?-24=2?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

What would V be?

OpenStudy (anonymous):

u try it first and tell me wat u get... the easiest thing would be to use all the answer choices to fill it in

OpenStudy (anonymous):

this one is hard..

OpenStudy (anonymous):

16?! :D

OpenStudy (anonymous):

go for it! :)